a) A phone company charges $19.00 per month plus $0.06 per call. Find a rational function that expresses the average cost (y) of a call in a month when (x) calls were made.
b)A phone company charges $20.20 per month plus $0.05 per call. How many calls can be made for an average cost of $0.20 per call.
I really just don't understand how to do it and I cannot find any good examples..so I would really appreciate if someone could help :-)
Thanks.

Answer by oscargut(686)


a) the cost is 0.06x + 19 = (6x+1900)/100 then y =(6x+1900)/100x
Answer: y = (6x+1900)/100x
b) cost = (0.05x+20.2) then average cost is (0.05x+20.2)/x
then (0.05x+20.2)/x = 0.2
then (0.05x+20.2) = 0.2x
then 0.15x=20.2
then x=20.2/0.15= 134,6666...
Answer: 135 calls
Are you sure about 20.20 and 0.05? (because solution is not an integer)